OpenStudy (anonymous):

What is the equation of the lines that passes through the poins of (0,1) and (6,-2)

Elsa213 (elsa213):

First you do need to find the slope of the line based on m=(y2-y1)/(x2-x1) so for these points: (0,3) and (2,0) --> (x1, y1) and (x2, y2) x1=0 y1=3 x2=2 y2=0 m= (0-3)/(2-0) = -3/2 Then your formula will need to look like y=mx+b where b is your y-intercept (when x=0) Due to have the x=0 point, you are pretty much done: y=-3/2*x+3 If you do not have a (0,y) form, you can solve as follows: Write out the formula with your slope value inserted: y=-3/2*x+b Choose one of your points on your line to solve for b. I choose (2,0): 0=-3/2*2+b 0=-3*1+b 3=b --> b=3

OpenStudy (anonymous):

Ok so you have two points given, so first lets find the slope. (0,1)(6,-2) is also (X1,Y1)(X2,Y2).